If the sum of first m terms of AP is n and sum of n terms of AP is m, then show that the sum of (m + n) terms is  "-(m+n)".

Grade:10

2 Answers

jitender lakhanpal
12 years ago

Dear sathya ,

 m = n/2[2a + (n-1)d]                                where a and d are the first term and the commom diffrerence 

 n = m/2[2a + (m-1)d]

2a + (n-1)d = 2m/n-----1

2a + (n-1)d = 2n/m --------2

now solving the two equations for the 2 values of a and d 

and then put the values of a and d in the equation

sum = (m+n)/2[2a + (m+n)d]

we will find that is equal to -(m+n)

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

YOUR NAME 

Sathya
12 years ago

?In the 4th step,isnt it 2a + (m-1)d = 2n/m ? If u solve this equations,i dont think we might be gettin the ans.Please show the steps.